OpenStudy (solomonzelman):

I suck at this... just need someone to check. It looks stupid, I am trying some problems online, and didn't even't learn this in school.

OpenStudy (solomonzelman):

§ cos^2 x sin(x) dx = (u=cos(x) ) § u^2sin(x) ( 1/-sin(x ) du = § -u^2 du = taking the constant out,v -§ u^2 du = applying the power rule -§ (-u^2+1) / (2+1) = sub in cos(x) for u ... = -cos{2+1}x / 2+1 = -cos^3x/3 My final answer is (-cos^3x/3)+C

OpenStudy (solomonzelman):

§ is supposed to be the (indefinite) integral

OpenStudy (solomonzelman):

reposting a little. I forgot a caret -:( sub in cos(x) for u ... = -cos^{2+1}x / 2+1 = -cos^3x/3 My final answer is (-cos^3x/3)+C

OpenStudy (luigi0210):

I think it might be easier to change that \(cos^2 x\) into \( 1-sin^2 x\)

OpenStudy (luigi0210):

Actually, no, sorry, you did it right :P

OpenStudy (luigi0210):

\[\Large \int cos^2 x ~sinx~ dx\] \(u=cosx\) \(du=-sinxdx\) \[\Large -\int u^2du\] \[\Large -\frac{u^3}{3} +C\] Sub back in~ \[\Large -\frac{cos^3~x}{3}+C\]

OpenStudy (solomonzelman):

Thank you for your clarification, Luigi !